HOT-133: Chief Scientist Report
Chief Scientist: T. GREGORY
Departed: December 12, 2001 at 1000 (HST)
Returned: December 16, 2001 at 1930 (HST)
Vessel: R/V Ka'imikai-o-Kanaloa
Operator: University of Hawaii
Master of the Vessel: Captain Robert Hayes
Chief Scientist: Thomas Gregory
STAG Electronics Technician: Steve Poulos
STAG Deck Operations: Dave Gravatt
1. SCIENTIFIC OBJECTIVES
The objective of this cruise was to continue building a collection of
hydrographic and biogeochemical data at the Hawaii Ocean Time-series
(HOT) stations. Four stations were to be occupied during the cruise, in
the following order:
1) Station 1, referred to as Station Kahe, is located at 21° 20.6'N, 158
° 16.4'W and was to be occupied on December 12 for about 3 hours.
2) Station 2: ALOHA (A Long Term Oligotrophic Habitat Assessment) is
defined as a circle with a 6 nautical mile radius centered at 22° 45'N,
158°W. This is the main HOT station and was to be occupied for 4 days
from December 13 to December 16.
3) Station 8, referred to as HALE-ALOHA is the location of our deep
ocean mooring (20° 20'N, 158° 10.6'W). The mooring is no longer
deployed. It was to be occupied on December 16 for about 2 hours.
4) Station 6, referred to as Station Kaena, is located off Kaena Point
at 21° 50.8'N, 158° 21.8'W was to be occupied on December 16 for about
A single CTD cast was to be conducted at Station 1 to collect
continuous profiles of various physical and chemical parameters. Water
samples were to be collected at discrete depths for biogeochemical
measurements. PRR and TSRB measurements were also to be made.
Upon arrival at Station ALOHA, net tows were to be conducted followed
by the deployment of a free-drifting sediment trap array. After
deployment, a full-depth CTD cast was to be conducted followed by CTD
casts at strict 3- hour intervals for at least 36 hours for continuous
and discrete data collection followed by another full-depth CTD cast.
Two other free-drifting arrays were to be deployed on December 14: an
oxygen balance experiment for 24 hours and a primary production
experiment for 12 hours. Plankton net tows were to be conducted near
noon and midnight on December 13 and 14 at Station ALOHA. PRR and TSRB
deployments and recoveries were to be done around noon December 13 and
14. After work at Station ALOHA was accomplished, the ship was to
transit to recover bottom moored sediment traps, which had been
deployed near Station ALOHA on December 15, 2000. Following this
operation we were to recover the drifting sediment trap array.
Following Station ALOHA operations, the ship was to transit to Station
8 to conduct one CTD cast and then transit to Station 6.
A near-bottom CTD cast (~2500 m) was to be conducted at Station 6
including salinity samples for calibration, after which the ship was to
transit back to Snug Harbor.
The following instruments were to collect data throughout the cruise:
a shipboard ADCP, a thermosalinograph and fluorometer, and an
2. SCIENCE PERSONNEL
Noel Larson (Watch Leader) Research Associate UH
Jeremiah Johnson Research Associate UH
Mark Valenciano Electronics Technician UH
Fernando Santiago-Mandujano Research Associate UH
Tom Gregory (Chief Scientist) Research Associate UH
Karin Björkman Research Specialist UH
Anne Gasc Research Associate UH
Lance Fujieki (Watch Leader) Computer Specialist UH
Paul Morris Technician UH
John Dore Scientist UH
Tara Clemente Research Associate UH
Colleen Allen Research Associate UH
3. GENERAL SUMMARY
One 1000-m CTD cast was obtained at both Station Kahe and Station
HALE-ALOHA however we did not complete a cast at Station Kaena.
PRR/TSRB deployments were successful at Station HALE-ALOHA. Heavy seas
and high winds hampered operations at Station ALOHA . After waiting
for the heavy weather to subside to operational levels, all CTD-related
sampling goals were accomplished including the 36-hour sampling
Neither free-floating array deployments nor PRR/TSRB work were
attempted at Station ALOHA due to extreme weather. Also, we decided to
postpone recovery of the bottom moored sediment trap to a future
C. Allen attempted one plankton net tow however most of the 202 ?m net
including the cod end was lost.
The ADCP ran without interruption throughout the cruise, as well as the
fluorometer, thermosalinograph and the ship's anemometer. However,
thermosalinograph conductivity data was quite noisy, probably due to
bubbles in the system.
We arrived back at Snug Harbor on December 16 at about 1930. A partial
off-load took place the next morning.
4. R/V KA'IMIKAI-O-KANALOA, OFFICERS AND CREW, TECHNICAL SUPPORT
The R/V Ka'imikai-o-Kanaloa and her crew continue to deliver excellent
ship support for our work. The officers and crew were most helpful and
accommodating and are to be commended for maintaining their high
standards during a grueling cruise. They showed enthusiasm and concern
for our work and were very flexible in receiving changes in our
operational schedule that were mandated by the extreme weather we
Technical support during this cruise was excellent. STAG personnel were
available at any time to assist in our work and made things much easier
5. DAILY REPORT OF ACTIVITIES (HST)
December 11, 2001; Loading Day
Equipment loaded on this day. Tested CTD system.
December 12, 2001
The ship departed from Snug harbor at 1015. Fire and abandon ship
drills were conducted at 1049, followed by a short science meeting
during which the cruise schedule was reviewed and safety issues were
We arrived at Kahe station at 1314 and the PRR and TSRB were deployed
immediately. Both worked without problems. A weight cast (400 lb) to
1000 m was conducted at 1350 followed by the CTD cast at 1506. A bottle
was leaking on retrieval and another didn't close. The STAG Seapoint
fluorometer (SN 2440) was used as the JGOFS fluorometer was not on
board. The CTD cast was back on board at 1610 and we began transit to
Station ALOHA. Due to high seas our speed was reduced to 7-8 knots.
Close to midnight the starboard forward vent box developed a leak.
Seawater was observed flooding into the starboard lounge at about 1
gal/min. Ship engineers placed a soft patch on the holed vent box and
we resumed transit to Station ALOHA at 2345. Sea conditions demanded a
reduced speed of 6 knots.
December 13, 2001
We arrived at station ALOHA at 0650. Sustained 35 knot winds and steep,
short period seas averaging 14 feet forced postponement of science
operations. Conditions worsened throughout the day precluding any
Thermosalinograph conductivity data was quite noisy. We believe there
were bubbles in the system due to high seas and extreme ship motion.
The ADCP showed a strong north current of about 1 knot turning
northwestward below 100 m.
December 14, 2001
News of forecasted weather abatement was received during a satellite
telephone with Roger Lukas, prompting us to remain on station. Heavy
weather continued throughout the day however conditions improved such
that we were able to begin our first CTD cast at 1950. A large
northwest swell was expected to arrive by around noon on December 15
and we were concerned that we had only a short window in which to do
some casts. Therefore we decided to begin with a tentative schedule of
six casts, including one deep cast, that addressed our core sampling
requirements. The first CTD cast was back on deck at 2055. The second
cast went in at 2225 and was recovered at 2328.
Marginal working conditions combined with strong surface currents
prompted the decision to cancel free-floating array and PRR/TSRB work
for this cruise.
A zooplankton net tow was begun at 2336. The net tow was recovered at
2358. The 202 ?m net had ripped away from the rig and was lost. No
more net tows were attempted this cruise.
December 15, 2001
The remaining four casts in our abbreviated schedule, including a deep
cast to approximately 4800 m, were successfully completed. At this
point all JGOFS goals involving CTD work had been met. We decided to
remain on station to complete the 36-hour sampling period required by
WOCE. We also decided to postpone recovery of our moored sediment trap
until another cruise.
A total of 8 CTD casts were successfully completed this day.
December 16, 2001
We successfully completed two CTD casts this morning, meeting the
36-hour sampling period goal. We began transit to Station 8 at 0550.
The ship arrived at Station 8 and we began the CTD cast at 0900. CTD
was recovered at 0953 and we departed for Snug Harbor.
The WOCE group decided to forego Station 6 this cruise.
Arrived at Snug Harbor at 1930.
December 17, 2001
A partial offload was completed this morning.
Below is the cruise bridge log description for HOT 133. Wind and sea
directions are in degrees, wind speed in knots, seas in Beaufort scale,
swells in feet, barometer in inches Hg, temp in ?F (dry bulb), clouds
Date Wind Sea Swell Barometer Temp Clouds
Wed 12, Dec. 080, 25-30 080, 5 120, 8-10 29.98-30.05 75-82 1-8
Thurs 13, Dec. 085-095, 32-38 085-095, 5-6 080-100, 15-20 30.04-30.10 72-78 4-9
Fri 14, Dec. 090-110, 26-40 090-110, 5-6 090, 16 29.99-30.09 73-76 6-10
Sat 15, Dec. 075-110, 17-26 075-110, 4-6 100, 10-15 29.96-30.05 74-76 4-10
Sun 16, Dec. 070-080, 20-24 070-080, 3-5 090-310, 5-10 30.00-30.05 71-73 2-7
Sub component programs:
Bob Bidigare HPLC pigments/UH
Mike Landry zooplankton dynamics/UH
John Dore CO2 dynamics/UH
Charles Keeling CO2 dynamics and intercalibration/SIO
Paul Quay DI13C and O isotopes/UW
Mark Abbott/Ricardo Letelier optical measurements/OSU
Peter J. LeB. Williams oxygen balance/U Wales Bangor, UK
Dale Hebel EOC production /UH
Karin Björkman phosphorus cycling/UH
Paul Morris oxygen balance/UH